Suppose that a 1984 the total output in a single good economy was 7000 buckets of chicken also suppose that 1984 each bucket of chicken was priced at $10 finally assume that in 2005 the price at per bucket of chicken with $16 that 22000 buckets of chickens were produced by what percentage did the price level as measured by this index rise between 1984 and 2005
Answer #1

Answer - 2008 is the base year

Price in base year = $10

Price in current year = $ 16

Quantity in base year = 18000

Cost of basket in base year = 18000*10

= $ 180000

Cost of same basket as per CY price = 18000*16

= $ 288000

Price index = 288000/180000*100

= 160 %

This means that the prices rose by 60 percent as compared to 2008